The properties of a closed system change following the relation between pressure and volume as p V = 3.0 where p is in bar V is in m3. Calculate the work done when the pressure increases from 1.5 bar to 7.5 bar.
1
Expert's answer
2019-03-05T09:44:25-0500

W=∫v1v2 pdV

V1 = 3.0/p1 = 3.0/1.5 = 2 m3

V2 = 3.0/p2 = 3.0/7.5 = 0.4 m3

W=∫20.4 (3.0/V) dV=3×[lnV]20.4=3×(ln0.4-ln2)= -4.83 barr×m3

As 1 barr = 105 N/m2, then W = -4.83*105 (N/m2)*m3 = -4.83*105 Nm = -4.83*105 J = -483 kJ

Answer: -483 kJ